# Object as a Service: Simplifying Cloud-Native Development through Serverless Object Abstraction

### v1

@@ VI-E Case Study: Development Efficiency Using OaaS
In this part, we provide a real-world use case of object development using OaaS and its FaaS counterpart and then demonstrate how OaaS makes the development process of cloud-native serverless applications easier and faster. The use case is a video processing application that performs face detection and facial expression recognition. Figure 11 shows the automatic system uploads the video file to the object storage to be processed by the workflow of functions. The workflow includes: Func_1 that splits the video into multiple image segments; Func_2 that detects the face on each sample image frame; and Func_3 performs the facial expression recognition on the detected face image and generates label data in the JSON format. These functions must persist their output object so that the next function in the workflow can consume it.
FaaS implementation. The developer must implement the following steps: (i) Configuring cloud-based object storage and managing access tokens. (ii) Implementing business logic to respond to trigger events. (iii) Manage data within the functions that involve three steps that are allocating the storage addresses, authenticating access to the object storage, and performing fetch and upload operations to the allocated addresses. Upon implementing these functions, the developer has to connect them as a workflow via a function orchestrator service. Finally, the dashboard service invokes the workflow upon receiving a user request and collects the results.
@@ VII Conclusions
In this research, we presented the OaaS paradigm that aimed at simplifying state and workflow management for cloud-native applications. Our prototype, Oparaca, supports both structured and unstructured data types, ensuring consistency through fail-safe state transitions. It also provides secure and low-overhead management for unstructured data using presigned URLs and redirection mechanisms. For structured data, it employs the pure function scheme to transparently manage application data to the developer code and using DHT and consistent hashing to scalably cache the object data and improve the data locality. To make the Oparaca fault-tolerant, we developed the exactly-once and localized locking schemes. To support cloud-native workflow, Oparaca enables declarative dataflow abstraction that hides the concurrency and synchronization concerns from the developer’s perspective. The evaluation results demonstrate that Oparaca streamlines cloud-native programming and is ideal for use cases that require persisting the state or defining a workflow. Oparaca offers scalability with negligible overhead, particularly for compute-intensive tasks. In the future, we plan to develop Oparaca to support application deployment across multiple data centers, streamlining large-scale application development.

### v2

@@ VI-E Case Study: Development Efficiency Using OaaS
In this part, we provide a real-world use case of object development using OaaS and its FaaS counterpart and then demonstrate how OaaS makes the development process of cloud-native serverless applications easier and faster.
Case Study # 1. expression detection system. This case study is a video processing workflow that performs face detection and facial expression recognition. Figure 11 shows the automatic system uploads the video file to the object storage to be processed by the workflow of functions. The workflow includes: Func_1 to split the video into multiple image segments; Func_2 to detect the face on each sample image frame; and Func_3 to perform facial expression recognition on the detected face image and generate a JSON format label. These functions must persist their output object so that the next function in the workflow can consume it.
FaaS implementation. The developer must implement the following steps: (a) Configuring cloud-based object storage and managing access tokens. (b) Implementing business logic to respond to trigger events. (c) Manage data within the functions that involve three steps: allocating the storage addresses, authenticating access to the object storage, and performing fetch and upload operations to the allocated addresses. Upon implementing these functions, the developer has to connect them as a workflow via a function orchestrator service. Finally, the dashboard service invokes the workflow upon receiving a user request and collects the results.
@@
(a) FaaS-based
(b) OaaS-based
Figure 12: The automatic content moderation system.
Case Study # 2. content moderation system. Moderating the content at scale in various formats, including image, document, and video. We first present how the application is deployed in FaaS, the limitations of this approach, and how OaaS can resolve those limitations.
FaaS implementation. To simplify complex multimedia processing workflows [8], we abstract the workflow to extract the metadata from the image files as Image workflow, and the workflow to extract metadata from the text files as Text workflow as shown in Figure 12a. Before using both workflows, the content must be pre-processed to extract raw data via using the pertinent FaaS functions: (a) text extraction to extract text from the document. (b) transcribing to extract text from the video. (c) frame extraction to sample image frame from the video. After feeding the data into both workflows to extract the metadata, we use the evaluation function to generate a report to the content moderation system.
The FaaS implementation has three main drawbacks: (A) developers must explicitly manage application state and data using separate storage services that increase the complexity. (B) even though the common workflow can be reused, the intermediate data management is not abstracted. That is, if the developer needs to separate/change the staging storage, it must be done manually. (C) functionalities may require numerous and heterogeneous FaaS deployments—e.g., requiring a separate workflow for each content type, where the Video format requires a more complicated workflow than the other types. These drawbacks complicate development, deployment, and management as the application evolves to handle various document types and integrates more functionalities and options (e.g., using multiple evaluation services instead of one).
OaaS implementation. To demonstrate the efficacy of OaaS, we transform the given FaaS-based solution into OaaS with minimal effort to resolve the aforementioned drawbacks.
  • Workflow Construct. We encapsulate related FaaS functions and states into classes representing two key functionalities: Media to extract the metadata and Evaluator to evaluate metadata and report to the content moderation system. The two classes form the critical path of the application processing pipeline, as shown in Figure 12b.
  • Object Encapsulation. We use inheritance and polymorphism to enhance software reuse by encapsulating FaaS functions and states in classes derived from two base classes. This approach hides the need for storage services behind object abstraction and allows their implementation to be managed in the cloud. It simplifies development, as developers only build the processing pipeline once in the base classes. They can then focus on specific functionalities in the derived classes, avoiding redundant implementation when adding new content types or evaluator services.
@@ VII Conclusions
In this research, we presented the OaaS paradigm, which aims to simplify state and workflow management for cloud-native applications. Our prototype, Oparaca, supports both structured and unstructured data types, ensuring consistency through fail-safe state transitions. It also provides secure and low-overhead management for unstructured data using presigned URLs and redirection mechanisms. For structured data, it employs the pure function scheme to transparently manage application data to the developer code and using DHT and consistent hashing to scalably cache the object data and improve the data locality. To make the Oparaca fault-tolerant, we developed the exactly-once and localized locking schemes. To support cloud-native workflow, Oparaca enables declarative dataflow abstraction that hides the concurrency and synchronization concerns from the developer’s perspective. The evaluation results demonstrate that Oparaca streamlines cloud-native programming and is ideal for use cases that require persisting the state or defining a workflow. Oparaca offers scalability with negligible overhead, particularly for compute-intensive tasks. In the future, we plan to develop Oparaca to support application deployment across multiple data centers, streamlining large-scale application development.
ACKNOWLEDGEMENT
We would like to thank anonymous reviewers for their constructive feedback; and Chameleon Cloud for providing resources. This project is supported by National Science Foundation (NSF) through CNS CAREER Award# 2419588.